OpenStudy (anonymous):

Did I do this Geometry problem correctly? Write the equation for the circle with center (-5, 6) and diameter= 16 Is it (x-5)^2+(y+6)^2=256?

OpenStudy (anonymous):

Someone is answering, finally! :D

OpenStudy (anonymous):

equation of circle (x-h)^2+(y-k)^2=r^2 here h=-5 and k=6 (x-(-5))^2+(y-6)^2=(16)^2

OpenStudy (anonymous):

you did it wrong

OpenStudy (sirm3d):

it should be the radius that is squared, not the diameter

OpenStudy (anonymous):

sorry i think radius is given so put r=16/2=8

OpenStudy (anonymous):

so it should be (x+5)^2 + (y-6)^2 = 256

OpenStudy (anonymous):

*(x+5)^2 + (y-6)^2 = 64?

OpenStudy (anonymous):

yes now u did id?

OpenStudy (sirm3d):

@BagelsAreIcelandic the right side is 64